OpenStudy (brucebaner):

1. (1 pt) Relationship B has a greater rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose all answers that are correct.

OpenStudy (brucebaner):

A y =3/4x B y =0.6 C Y=2/3x D Y=1/4x

OpenStudy (ccswims):

can you post the graph?

OpenStudy (brucebaner):

@ccswims

OpenStudy (brucebaner):

@Destinymasha

OpenStudy (ccswims):

Well equation A has a slope of 1/2. So what ever equations listed above that have a slope of less than 1/2 are correct

OpenStudy (ccswims):

for example, A you can cross off right away because it has a a bigger slope than 1/2

OpenStudy (ccswims):

OH WAIT IM SORRY, I didn't read the question right, Equation B is greater than Equation A

OpenStudy (ccswims):

Ok, So A is actually one that you would circle because it has a greater slope of 1/2

OpenStudy (ccswims):

C is also correct because it has a greater slope of 1/2

OpenStudy (ccswims):

So A and C are both correct